# Causal machine learning methods and use of sample splitting in settings with high-dimensional confounding

### v1

@@ 9 Discussion
The simulation study was informed by a realistic case study, where we attempted to simulate confounders with dependencies consistent with associations observed between them in our realistic dataset. We considered several scenarios, representing both simple and complex data generating mechanisms, and simultaneously evaluated performance of the methods across a range of modest sample sizes that are likely to be encountered in observational studies. However, although we considered a wide range of scenarios, further data generating mechanisms could be considered to control the heterogeneity of the ACE across scenarios. Indeed, in the complex scenarios, for the outcome generation models, while the main effect coefficient value for the exposure was modified depending on sample size to control power (and thus, ensure comparability of coverage probability estimates across scenarios), coefficients for the interaction terms were not modified depending on sample size. This could have induced greater heterogeneity in the causal effect across strata in some scenarios, making it more challenging to compare bias of point and SE estimates across scenarios. In this work, the exposure was dichotomised. In addition to the threshold being arbitrary, as mentioned, we acknowledge that there are many interventions that may bring GlycA concentration below the 75th percentile and each of these could have different effects on cardiovascular health, that is, the intervention remains ill-defined. This is an inevitable feature of discovery phase studies of this kind. Further aspects of SL implementation that could potentially affect CML method performance (Bach et al., 2024) were not fully explored. For example, we did not consider the tuning of hyperparameters for the learners and did not compare the use of screening vs no screening in SL implementations as we only considered libraries without separate explicit screening. The simulation was based on one realistic motivating example, which may limit the guidance that is able to be provided from this study.
@@
Future research directions include exploring the impact of the sample splitting variant (e.g., double CF compared to single CF) on performance of the methods. Double CF may possibly result in faster convergence rates compared to single CF (Meng and Huang, 2022). However, Zivich and Breskin (2021) discuss the potential challenge of using double CF with small sample sizes and suggest that single CF may overcome some of the difficulty. Limited sample sizes are common in realistic settings, and it remains for future research to determine whether double CF may be preferable to single CF in these situations. We have shown that further research is required into the composition of SL libraries, particularly as there is a need to balance the requirement of sufficient diversity of learners with performance across an extensive range of data-generating mechanisms, with or without sparsity. It would also be helpful to explore any potential benefits of applying explicit screening within SL itself in a high-dimensional confounding setting. Further possibilities include extending this work beyond dichotomous exposures (e.g. continuous exposures) as they are often of interest, as well as examining challenges posed by missing data. In the high-dimensional confounding setting, evaluation, and development of the CML methods in the presence of missing data is important. This is a growing area of research (Dashti et al., 2024).
@@
To conclude, in a high-dimensional confounding setting, we found that on most occasions AIPW and TMLE performed similarly for estimating the ACE, with a slight preference toward the use of TMLE, given its better stability. In cases where complex mechanisms are suspected, a more diverse SL library that includes both parametric and non-parametric learners may be beneficial to reduce bias in point estimates and variance. CF is important for the estimation of SE and should be used, as without it we observed underestimation of model-SE and undercoverage, particularly with a more diverse library. Our study suggests that in the high-dimensional confounding setting, applying the methods with CF is most important, with the number of folds used with CF less important. Using many folds in CF may add computational burden, without substantial benefit to method performance.
10 Software
R code, together with a simulated input data set and complete documentation is available on request from the corresponding author.
Supplementary material is available online.
Acknowledgments
@@
We thank the BIS investigator group for providing access to the case-study data for illustrative purposes in this work. In addition, the authors thank David Burgner and Toby Mansell for sharing their expertise in the growing cardiovascular research area.
Funding
@@
This work was supported by the Australian National Health and Medical Research Council (NHMRC) Investigator Grant Emerging Leadership Level 2 (grant 2009572 awarded to MM-B). Research at the Murdoch Children’s Research Institute is supported by the Victorian Government’s Operational Infrastructure Support Program. This research was also supported by an Australian Government Research Training Program (RTP) Scholarship.
Conflict of Interest: None declared.

### v2

@@ 9 Discussion
Strengths of our simulation study include that it was informed by a realistic case study, where we attempted to simulate confounders with dependencies consistent with associations observed between them in our realistic dataset. We considered several scenarios, representing both simple and complex data generating mechanisms, and simultaneously evaluated performance of the methods across a range of modest sample sizes that are likely to be encountered in observational studies. However, although we considered a wide range of scenarios, further data generating mechanisms could be considered to control the heterogeneity of the ACE across scenarios. Indeed, in the complex scenarios, for the outcome generation models, while the main effect coefficient value for the exposure was modified depending on sample size to control power (and thus, ensure comparability of coverage probability estimates across scenarios), coefficients for the interaction terms were not modified depending on sample size. This could have induced greater heterogeneity in the causal effect across strata in some scenarios, making it more challenging to compare bias of point and SE estimates across scenarios. Further aspects of SL implementation that could potentially affect CML method performance 56 were not fully explored. For example, we did not consider the tuning of hyperparameters for the learners and did not compare the use of screening vs no screening in SL implementations as we only considered libraries without separate explicit screening. A further limitation of our study is that we did not examine the method for cross-fitting that aggregates over different random seeds to do the split, suggested by Chernozhukov et al 23, which could be more robust than the methods we assessed that relied on a single seed/split. Finally, the simulation was based on one realistic motivating example, which may limit the guidance that is able to be provided from this study.
@@
Future research directions could explore the CML methods in applications with a rare treatment or a binary, possibly rare outcome, or a large number of categorical confounders, all of which could bring about challenges with the application of sample splitting. Exploring the impact of the sample splitting variant (e.g., double CF compared to single CF) on performance of the methods would also provide insight. Double CF may possibly result in faster convergence rates compared to single CF. 26 However, Zivich and Breskin 3 discuss the potential challenge of using double CF with small sample sizes and suggest that single CF may overcome some of the difficulty. Limited sample sizes are common in realistic settings, and it remains for future research to determine whether double CF may be preferable to single CF in these situations. We have shown that further research is required into the composition of SL libraries, particularly as there is a need to balance the requirement of sufficient diversity of learners with performance across an extensive range of data-generating mechanisms, with or without sparsity. It would also be helpful to explore any potential benefits of applying explicit screening within SL itself in a high-dimensional confounding setting. Further possibilities include extending this work beyond dichotomous exposures (e.g. continuous exposures) as they are often of interest. Possibilities for continuous exposures include exploring a shifting exposure distribution 57, 58, 59, 60 or estimation of standard regression parameters via an influence curve approach 61. The latter approach focusses on a summary of the exposure effect with reduced modelling assumptions, rather than ascertaining how the outcome changes with step increases or changes in exposure. It will also be important to explore challenges posed by missing data. In the high-dimensional confounding setting, evaluation, and development of the CML methods in the presence of missing data is important. This is a growing area of research. 62
@@
To conclude, in a high-dimensional confounding setting, we found that on most occasions AIPW and TMLE performed similarly for estimating the ACE, with a slight preference toward the use of TMLE, given its better stability. In cases where complex mechanisms are suspected, a more diverse and flexible SL library may be beneficial to reduce bias in point estimates and variance. CF is important for the estimation of SE and should be used, as without it we observed underestimation of model-SE and undercoverage, particularly with a more diverse library. Our study suggests that in the high-dimensional confounding setting, applying the methods with CF is most important, with the number of folds used with CF less important. Using many folds in CF may add computational burden, without substantial benefit to method performance.
[…]
*Author contributions
All authors contributed to the planning and design of the simulation study. SE conducted the simulation study and performed the analyses, and prepared the first draft of the manuscript. JC, SV and MM-B reviewed the manuscript, and SE revised the manuscript accordingly. All authors reviewed, sighted and approved the final version of the manuscript.
[…]
@@
*Acknowledgments We thank the BIS investigator group for providing access to the case-study data for illustrative purposes in this work. In addition, the authors thank David Burgner and Toby Mansell for sharing their expertise in the growing cardiovascular research area.
[…]
@@
*Financial disclosure This work was supported by the Australian National Health and Medical Research Council (NHMRC) Investigator Grant Emerging Leadership Level 2 (grant 2009572 awarded to MM-B). SE is funded by an Australian Government Research Training Program Scholarship. Research at the Murdoch Children’s Research Institute is supported by the Victorian Government’s Operational Infrastructure Support Program.
[…]
*Conflict of interest
The authors declare no potential conflict of interests.